OpenStudy (anonymous):

7w - 3(4w + 8) = 11 Solve for w?

OpenStudy (anonymous):

7w-12w-24=11 -5w=35 w= -7

OpenStudy (dangerousjesse):

Distribute -3 over \(4 w+8.\) \(-3 (4 w+8) = -24-12 w:\) \(7 w+-24-12 w = 11\) Group like terms in \(7 w-12 w-24.\) Grouping like terms, 7 w-12 w-24 = (7 w-12 w)-24: \((7 w-12 w)-24 = 11\) Combine like terms in \( 7 w-12 w.\) \(7 w-12 w = -5 w:\) \(-5 w-24 = 11\) Isolate terms with w to the left hand side. Add 24 to both sides: \((24-24)-5 w = 11+24\) Look for two terms that sum to zero. \(24-24 = 0:\) \(-5 w = 11+24\) Evaluate \(11+24.\) \(11+24 = 35:\) \(-5 w = 35\) Divide both sides by a constant to simplify the equation. Divide both sides of \(-5 w = 35\) by -5: \(\frac{-5 w}{-5} = \frac{35}{-5}\) Any nonzero number divided by itself is one. \(\frac{-5}{-5} = 1:\) \(w = \frac{35}{-5}\) Reduce \(\frac{35}{-5}\) to lowest terms. Start by finding the GCD of 35 and -5. The gcd of 35 and -5 is 5, so \(\frac{35}{-5} = \frac{5×7}{5 (-1)} = \frac{5}{5}×\frac{7}{-1} = \frac{7}{-1}:\) \(w = \frac{7}{-1}\) Simplify the sign of \(\frac{7}{-1}\). Multiply numerator and denominator of \(\frac{7}{-1}\) by -1

OpenStudy (dangerousjesse):

Sorry for being so slow :P

OpenStudy (anonymous):

WOAHHHH nice!!!